Compare and contrast properties of waves and particles, explaining which properties can and cannot distinguish between wave-like behaviour and particle-like behaviour.
1
Expert's answer
2020-02-04T09:21:46-0500

Properties of waves: wavelength "\lambda", frequency f and velocity v. Properties of particles: mass m, charge q, and other quantum numbers. Mass and wavelength, according to De Broglie wavelength equation, can help to distinguish what it is - a wave or a particle:


"\lambda=\frac{h}{mv}."

The shorter the wavelength, the more the object behaves like a particle. On the other hand, this can be also said in terms of frequency:


"\lambda=\frac{v}{f}=\frac{h}{mv},\\
\space\\
f=\frac{mv^2}{h}."

Higher frequencies define particle-like behavior.

Waves can diffract and interfere, while particles cannot. We can say that it is possible to apply Newton's mechanics and kinematics to particles, but not to waves.